OpenStudy (ttop0816):

Solve 3cos2θ + 5cosθ + 2 = 0 for 0° ≤ θ < 180°

OpenStudy (anonymous):

Do you know the trigonometric identity for double angles? You should begin by changing cos2θ so that you have only cosθ to work with.

OpenStudy (ttop0816):

Ø {109°} <<< this one? {132°}

OpenStudy (anonymous):

Cos2θ = cos^2θ - sin^2θ You can change this so that you are only dealing with cosθ.

OpenStudy (ttop0816):

how would i go with the equation there??

OpenStudy (anonymous):

Do you know any identities? You must if you've been given this question. sin^2θ = 1-cos^2θ should help.

OpenStudy (tkhunny):

Is it really \(\cos(2\theta)\) or did you intend \(\cos^{2}(\theta)\)?

OpenStudy (ttop0816):

??

OpenStudy (anonymous):

lol latex fail days now

OpenStudy (anonymous):

the question is, is it cosine of two x, or cosine squared of x cos(2x) or cos^2(x)?

OpenStudy (anonymous):

i am going to guess it is 3cos^2(x)+5cos(x)+2=0 because this one factors as (cos(x) + 1)(3cos(x) + 2)=0 that gives you cos(x) = -1, or cos(x) = -2/3

OpenStudy (anonymous):

cos(x) = -1 tells you x = 180

OpenStudy (anonymous):

cos(x) = -2/3 you need a calculator to solve find arccos(-2/3)

OpenStudy (ttop0816):

132 degrees!

OpenStudy (anonymous):

@KinzaN fyi "latex" is the math writing tool here, and it is not working for the last few days that is why we have to write by hand -2/3 instead of \(-\frac{2}{3}\)

OpenStudy (anonymous):

yes, 132 rounded

OpenStudy (tkhunny):

I keep hearing about LaTeX failures, but I have yet to see one on this round. I am using FireFox. Maybe it's an IE thing?